(b) Given f(x) = x
2
, g(x) = 2x and h(x) = x - 2 be functions from R to R. Find
(i) f â—¦ g
(ii) f â—¦ h
(iii) h â—¦ f
(iv) (f â—¦ g) â—¦ h
(v) f â—¦ (g â—¦ h)
1
Expert's answer
2020-08-17T18:32:01-0400

(i)

"f\circ g=f(g(x))=f(2x)=(2x)^2=4x^2"

(ii)

"f\circ h=f(h(x))=f(x-2)=(x-2)^2=x^2-4x+4"

(iii)

"h\circ f=h(f(x))=h(x^2)=x^2-2"

(iv)

"(f\circ g)\circ h=(f\circ g)(h(x))=(f\circ g)(x-2)="


"=f(g(x-2))=f(2(x-2))=(2(x-2))^2="


"=4x^2-16x+16"

(v)


"f\circ (g\circ h)=f((g\circ h)(x))=f(g(h(x)))="

"=f(g(x-2))=f(2(x-2))=(2(x-2))^2="


"=4x^2-16x+16"